OpenStudy (anonymous):

The measures of the angles of a triangle are in the extended ratio 3:6:11. what is the measure of the largest angle?

OpenStudy (anonymous):

??

OpenStudy (mathstudent55):

The side that has length 11 is too long to work with sides of lengths 3 and 6 to be able to have a triangle.

OpenStudy (mathmale):

Yes: in other words, if the sides of length 3 and 6 were to be lined up straight, their combined length would come to 9, whereas the length of the longest side would be 11. No soap.

OpenStudy (anonymous):

well the question isnt wrong. idk=/ something like 20x +6x + 11x = 180 ??

OpenStudy (mathstudent55):

Sorry. Ignore what I wrote above. I misread the question.

OpenStudy (mathstudent55):

It's the angles that are in that ratio, not the sides. My mistake.

OpenStudy (anonymous):

the answer is 99 .. i think. thats my guess.

OpenStudy (mathmale):

Oh, shoot. RED face. We were talking about side lengths here; the problem deals with ANGLES, not side lengths. If x is a constant of proportionality, then 3x+6x+11x=180 degrees (because the interior angles of a triangle always sum up to 180 degrees).

OpenStudy (mathstudent55):

Since the angles are in the 3:6:11 ratio, let's call the angles 3x : 6x : 11x That means 3x + 6x + 11x = 180

OpenStudy (mathstudent55):

Now we solve for x: 3x + 6x + 11x = 180 20x = 180 x = 9

OpenStudy (mathmale):

Your guess? Just add up 3x and 6x and 11x and set them equal to 180; result is 20x=180. x+??

OpenStudy (anonymous):

9 isnt an option

OpenStudy (anonymous):

so 20?

OpenStudy (mathstudent55):

Since we know x = 9, and the angle is 11x, we replace the x with 9: 11x = 11*9 = 99

OpenStudy (anonymous):

yay i was right :D thanks!

OpenStudy (mathmale):

Right. 9 is nothing more than a constant multiplier. 9 is not a measure of one of the angles. Mathstudent has the right idea.